HYDROGEN BOND NETWORK ISOMERS OF THE WATER NONAMER AND DECAMER OBSERVED BY BROADBAND ROTATIONAL SPECTROSCOPY

\underline{CRISTOBAL PEREZ}, DANIEL P. ZALESKI, NATHAN A. SEIFERT, BROOKS H. PATE, Department of Chemistry, University of Virginia, McCormick Rd., Charlottesville, VA 22904-4319; ZBIGNIEW KISIEL, Institute of Physics, Polish Academy of Sciences, Al. Lotników 32/46, 02-668 Warszawa, Poland; BERHANE TEMELSO, GEORGE C. SHIELDS, Dean’s Office, College of Arts and Sciences, and Department of Chemistry, Bucknell University, Lewisburg, PA 17837.

After our previous study of the rotational spectrum of water clusters in the 6-18 GHz^{a} region, in order to study clusters of larger size (>8 water molecules), a chirped-pulse Fourier transform microwave spectrometer in the 2-8 GHz frequency range has been used to obtain the broadband rotational spectra of five water nonamer isomers and four water decamer isomers in a pulsed molecular beam. The oxygen atom framework geometries for three nonamers and two decamers have also been unambiguously identified from isotopic labeling measurements using an H_{2}^{18}O enriched sample. Three of the four observed water decamer show tunneling effect associated with the internal dynamics of hydrogen-bond network in a similar fashion as the prism water hexamer. These tunneling paths are quenched upon a single incorporation of a H_{2}^{18}O molecule in the cluster. Due the large amount of closely-spaced rotational transitions in the H_{2}^{18}O spectrum, automated fitting tools were employed to extract the corresponding rotational spectra, which will be also briefly described.

aC. Perez, M. T. Muckle, D. P. Zaleski, N. A. Seifert, B. Temelso, G. C. Shields, Z. Kisiel, and B. H. Pate, Science 336, 897